Help me please!

An aging of a company's accounts receivable indicates that the estimate of uncollectible accounts totals $5,927. If Allowance for Doubtful Accounts has a $1,220 debit balance, the adjustment to record the bad debt expense for the period will require a

Select the correct answer.

a. debit to Bad Debt Expense for $5,927.

b. debit to Bad Debt Expense for $1,220.

c. credit to Allowance for Doubtful Accounts for $5,927.

d. debit to Bad Debt Expense for $7,147.

Allowance for Doubtful Accounts has a debit balance of $453 at the end of the year (before adjustment), and Bad Debt Expense is estimated at 3% of sales. If net sales are $940,100, the amount of the adjusting entry to record the estimate of the uncollectible accounts is

a. $28,656

b. $27,750

c. $453

d. $28,203

An asset was purchased for $128,000.00 on January 1, Year 1 and originally estimated to have a useful life of 8 years with a residual value of $9,500.00. At the beginning of the third year, it was determined that the remaining useful life of the asset was only 4 years with a residual value of $2,200.00. Calculate the third-year depreciation expense using the revised amounts and straight line method.

Select the correct answer.

$25,043.75

$24,043.75

$23,043.75

$24,543.75
Answer #1

Answers

  • Question #1

--Correct Answer = Option ‘D” Debit to Bad Debt expense for $ 7,147

Adjusted balance required in Allowance account

$5,927

Credit

Actual unadjusted balance of Allowance account

$1,220

Debit

Bad Debt Expense

$7,147

[$ 1220 for clearing debit balance PLUS $ 5927 for making balance credit]

  • Question #2

--Correct Answer = Option ‘d’ $ 28,203

Net Sales

$940,100

Bad Debt expense rate

3% of net sales

Amount of adjusting entry [940100 x 3%]

$28,203

  • Question #3

--Correct Answer = Option #2: $ 24,043.75

A

Cost

$           128,000.00

B

Residual Value

$                9,500.00

C=A - B

Depreciable base

$           118,500.00

D

Life [in years]

8

E=C/D

Annual SLM depreciation

$              14,812.50

F = Ex 2

2 Year accumulated depreciation

$              29,625.00

G = A - F

Book Value at the time of revision

$              98,375.00

H

Revised residual value

$                2,200.00

I = G - H

New Depreciable base

$              96,175.00

J

New remaining life

                                  4

K = I/J

New revised depreciation expense from 3rd year

$              24,043.75